Idk how to do these and need a lot of help because i have a whole worksheet. So can someone plz help me step-by-step one #4.
1. Through: (1,4), Slope=9
2. Through: (-1,-3), Slope=5
3. Through: (-3,1), Slope=-5/3
*4. Through: (2,-3), Slope=undefined
5. Through: (5,-3), Slope=-1/5
6. Through: (3,-2), Slope=1
7. Through: (2,5), Slope=3
8. Through: (-2,-2), Slope=-1
9. Through: (-5,3), Slope=-8/5
10. Through: (3,-1), Slope=2/3
11. Through: (2,-1), Slope=-3
12. Through: (-1,-1), Slope=-4

Answer by Alan3354(69443)

1. Through: (1,4), Slope=9
Use y - y1 = m*(x - x1) where (x1,y1) is the point and m = slope
=====================================
2. Through: (-1,-3), Slope=5
3. Through: (-3,1), Slope=-5/3
Same as #1
-----
*4. Through: (2,-3), Slope=undefined
Slope undefined means "vertical," or parallel to the y-axis.
The x value is constant.
--> x = -3
======================================
5. Through: (5,-3), Slope=-1/5
6. Through: (3,-2), Slope=1
7. Through: (2,5), Slope=3
8. Through: (-2,-2), Slope=-1
9. Through: (-5,3), Slope=-8/5
10. Through: (3,-1), Slope=2/3
11. Through: (2,-1), Slope=-3
12. Through: (-1,-1), Slope=-4
Same as #1